QUANTUM INTERFERENCE PHENOMENON IN THE COLD ATOMIC CASCADE SYSTEM

\underline{RAY-YUAN CHANG}, Department of Physics, National Cheng-Kung University, Tainan 70101,Taiwan(R.O.C.); WEI-CHIA FANG, Department of Physics, National Cheng-Kung University, Tainan 70101, Taiwan(R.O.C.); MING-DA TSAI, Department of Physics, National Cheng-Kung University, Tainan 70101, Taiwan(R.O.C.); CHIN-CHUN TSAI, Department of Physics, National Cheng-Kung University, Tainan 70101, Taiwan(R.O.C.).

By using the technique of electromagnetically induced transparency(EIT), we have observed the quantum interference phenomenon in the lower-lying atomic Rydberg states of cold Cs atom in a magneto-optical trap (MOT) with typical atom numbers up to 1×10^{7}, and cloud temperature about 100μK. In our experiments, a grating feedback diode laser is frequency stabilized on the Cs atom via Doppler-free saturation absorption transition |6s ^{2}S_{1/2}, F = 4 >→ |6p ^{2}P_{3/2}, F′ = 5 >, and a tunable Ti:Sapphire laser scans across the |6p ^{2}P_{3/2}, F′ = 5 >→ |8s ^{2}S_{1/2}, F′′ = 4 > transition frequency. Both spontaneous emission fluorescence from MOT and transmission intensity of the diode laser are monitored. A subnatural linewidth signal is obtained for probing the transition of Cs |6p ^{2}P_{3/2}, F′ = 5 >→ |8s ^{2}S_{1/2}, F′′ = 4 >. These signals correspond to the cascade EIT among the |6s ^{2}S_{1/2}, F = 4 >↔ |6p ^{2}P_{3/2}, F′ = 5 >↔ |8s ^{2}S_{1/2}, F′′ = 4 >.

This project is financially sponsored by National Science Council (grand no. NSC-95-2112-M-012-MY3).